OpenStudy (anonymous):

Which conjunction or disjunction is equivalent to the open sentence |a – 9| ≤ 4. A. a – 9 ≥ 4 and a – 9 ≤ –4 B. a – 9 ≤ 4 or a – 9 ≥ –4 C. a – 9 ≤ 4 and a – 9 ≤ –4 D. a – 9 ≤ 4 and a – 9 ≥ –4

OpenStudy (anonymous):

\[|a-9|\leq 4\\ -4\leq a-9\leq 4\]

OpenStudy (anonymous):

which his in fact D, not B

OpenStudy (anonymous):

it is an "and" statement, not an "or" statement